# SMC Is All You Need: Parallel Strong Scaling

### v1

@@ 2.1 Pre-conditioned Crank-Nicolson kernel
The scaling matrix $D$ should be chosen according to the information present in the likelihood. A simple and computationally convenient choice which is particularly amenable to use within SMC is to build it from an approximation of the target covariance [Beskos et al., 2018]. In the present work we adopt the simplest and cheapest choice and let $D=\text{diag}(\widehat{\text{var}}[\theta])$. The approximation of the variance, $\widehat{\text{var}}$, will be built from the current population of samples in the SMC case, and adaptively constructed in the MCMC case [Haario et al., 2001, Chen et al., 2016]. See also [Rudolf and Sprungk, 2018, Beskos, 2014, Zahm et al., 2022].
@@
  Give a current state $\theta$, the distribution $\pi_{j}$, a scaling parameter $\beta$.
@@
2.2 Hamiltonian Monte Carlo
@@
  Give a current state $\theta$, the distribution $\pi_{j}$ and a mass matrix $M_{0}$.
  Generate a initial momentum $p\sim N(0,M_{0})$.
@@
        Generate $\theta_{l\Delta t}$ and $p_{l\Delta t}$ from (6) with $\theta_{0}=\theta$ and $p_{0}=p$
@@
     if $u\leq\min\Big{\{}1,\frac{H(\theta^{\prime},p^{\prime})}{H(\theta,p)}\Big{\}}$ then
        $\theta=\theta^{\prime}$ and $p=p^{\prime}$.
@@
Hamiltonian Monte Carlo (HMC) kernel [Duane et al., 1987, Neal et al., 2011, Houlsby et al., 2011, Cobb and Jalaian, 2021] is essentially a gradient-based MCMC kernel on an extended state-space. We first build a Hamiltonian $H(\theta,p)$ with additional auxiliary “momentum” vector $p$ of the same dimension as the state $\theta$
$H(\theta,p)=-\log\pi_{j}(\theta)+\frac{1}{2}p^{\top}M_{0}^{-1}p,$ (4)
where $M_{0}$ is a mass matrix, so that $\frac{1}{Z}\exp(H(\theta,p))$ is a target distribution on the extended space, where the momentum can be simulated exactly $p\sim N(0,M_{0})$.
@@
$\frac{d\theta}{dt}=\frac{\partial H}{\partial p}=M_{0}^{-1}p,$ (5)
$\frac{dp}{dt}=\frac{\partial H}{\partial\theta}=\nabla_{\theta}\log\pi_{j}(\theta)\,.$
A carefully constructed symplectic-integrator is capable of conserving energy as well, for example the leapfrog integrator Leimkuhler and Reich [2004]:
$p_{t+\Delta t/2}=p_{t}+\frac{\Delta t}{2}\frac{dp}{dt}(\theta_{t}),$ (6)
$\theta_{t+\Delta t}=\theta_{t}+\Delta t\frac{d\theta}{dt}(p_{t+\Delta t/2}),$
$p_{t+\Delta t}=p_{t+\Delta t/2}+\frac{\Delta t}{2}\frac{dp}{dt}(\theta_{t+\Delta t}),$
@@
    simulate a random momentum $p\sim N(0,M_{0})$ (hence jumping to a new energy contour);
@@
    approximate the Hamiltonian dynamics using $L$ steps from (6);
@@
    correct numerical error from (ii) with an MH accept/reject step for $\frac{1}{Z}\exp(H(\theta,p))$.
@@
3 Parallel SMC method
By separating $NR$ samples into $R$ processors with $N$ samples in each, the parallel SMC sampler has a $R$ times lower time complexity than a single SMC sampler. This reduction in time complexity is crucial if we encounter an expensive SMC, which is common in the high-dimensional Bayesian inference problem. More importantly, under reasonable and mild assumptions we will demonstrate the strong scaling property for pSMC: the cost decreases linearly with the number of cores with constant time. The Algorithm 4 displays the parallel SMC method.
Algorithm 4 Parallel SMC sampler
  for $r=1$ to $R$ do
     Run Algorithm 1, output $\theta^{N,r}$.
     Compute the estimators $Z^{N,r}$, $f^{N,r}(\varphi)$, $f^{N,r}(1)$ and $\pi^{N,r}(\varphi)$ by (2).
  end for
$F^{N,R}(\varphi)$ $=\frac{1}{R}\sum_{r=1}^{R}f^{N,r}(\varphi),$ (7)
$F^{N,R}(1)$ $=\frac{1}{R}\sum_{r=1}^{R}f^{N,r}(1)=\frac{1}{R}\sum_{r=1}^{R}Z^{N,r}.$ (8)
$\hat{\varphi}_{\text{pSMC}}=\frac{F^{N,R}(\varphi)}{F^{N,R}(1)}=\frac{\sum_{r=1}^{R}f^{N,r}(\varphi)}{\sum_{r=1}^{R}f^{N,r}(1)}.$ (9)
3.1 Theoretical results for parallel strong scaling
In order to make use of (9), we define $f^{N,r}(\zeta)$ both for quantity of interest $\zeta=\varphi$ and $\zeta=1$. We denote that $|\zeta|_{\infty}=\max_{\theta\in\mathsf{\Theta}}|\zeta(\theta)|$ in the following equations. $C_{\zeta}$ denotes the constant depended on the function $\zeta$. Note that using $\varphi$ with one-dimensional output is without loss of generality for our convergence results, and the following proof can be directly generalized to the multi-output function by using the inner product.
Assumption 3.1.
@@ 3.1 Theoretical results for parallel strong scaling
Assumption 3.2.
Let $J\in\mathbb{N}$ be given, for each $j\in\{1,...,J\}$, there exists a $\rho\in(0,1)$ such that for any $(u,v)\in\mathsf{\Theta}^{2}$, $A$ is a measureable set from the space containing all the measureable subset of $\mathsf{\Theta}$,
@@
First, we restate the convergence result for the non-parallelized SMC method in Proposition 3.3, which is from Del Moral [2004].
Proposition 3.3.
Assume Assumption 3.1 and 3.2. Then, for any $J\in\mathbb{N}$, there exists a $C>0$ such that for any $N\in\mathbb{N}$, suitable $\zeta:\mathsf{\Theta}\rightarrow\mathbb{R}$,
@@
Lemma 3.4.
@@
Assume Assumption 3.1 and 3.2. Then, for any $J\in\mathbb{N}$, there is a $C>0$ such that for any $N,R\in\mathbb{N}$, suitable $\zeta:\mathsf{\Theta}\rightarrow\mathbb{R}$,
$\mathbb{E}[(F^{N,R}(\zeta)-f(\zeta))^{2}]\leq\frac{C|\zeta|_{\infty}^{2}}{NR}.$
Theorem 3.5.
@@
Assume Assumption 3.1 and 3.2. Then, for any $J\in\mathbb{N}$, there exists a $C_{\varphi}>0$, which depends on $\varphi$, such that for any $R\in\mathbb{N}$,
[…]
Then the strong scaling in the parallel SMC is described as the MSE defined in (11) decreases as $1/R$ for $R$ nodes, for suitable variables value ($M$,$N$,$J$), with constant time.
3.2 Discussion of pSMC in practice
Table 1: Comparison between pCN and pSMC regarding to the time cost, memory cost and MSE.
Time Cost Memory Cost MSE
pCN COST$(d,m)NR$ $d(m+NR)$ $C_{\sf MCMC}^{2}(d,m)/NR$
pSMC ${\sf COST}(d,m){\sf COST}_{\sf over}(d,m)N$ $d(m+N)$ $C_{\sf SMC}^{2}(d,m)/NR$
Let $m$ denote the number of observations in the data. As Theorem 3.5 suggests, the nuisance parameters $(M,N,J)$ must be chosen large enough to ensure rapid convergence in $R$. For MCMC the error will be bounded by $C_{\sf MCMC}(d,m)/NR$, and the time (and total) cost will grow like ${\sf COST}(d,m)NR$, where ${\sf COST}$ is the basic cost of evaluating the likelihood (target)  [^2] We are simplifying discussion by referring only the the major sources of complexity $(d,m)$, while any other metric for problem complexity will also come into play. For example, under partial observations the linear model will become more difficult as the variance of the observations $\sigma^{2}$ tends to 0.. Meanwhile, the error of SMC will be bounded by $C_{\sf SMC}(d,m)/N$, while the time cost will grow like ${\sf COST}(d,m)MJN$. Since $M$ and $J$ will also depend on $(d,m)$ – and need to be chosen larger for more complex targets – we will denote by ${\sf COST}_{\sf over}(d,m):=MJ$ the cost overhead of SMC vs MCMC for a given total number of samples.
Finally, we can achieve a fixed MSE of $\varepsilon^{2}$ with MCMC by selecting $NR=C_{\sf MCMC}(d,m)\varepsilon^{-2}$, which incurs a complexity of
${\sf COST}(d,m)NR={\sf COST}(d,m)C_{\sf MCMC}(d,m)\varepsilon^{-2}\,.$
Similarly, the time complexity of SMC will be
${\sf COST}(d,m){\sf COST}_{\sf over}(d,m)N={\sf COST}(d,m){\sf COST}_{\sf over}(d,m)C_{\sf SMC}(d,m)\varepsilon^{-2}/R\,.$
Equating these two, we can estimate the crossover
$R>{\sf COST}_{\sf over}(d,m)C_{\sf SMC}(d,m)/C_{\sf MCMC}(d,m)\,.$
For suitably chosen $M,J$, we expect $C_{\sf SMC}(d,m)\approx C_{\sf MCMC}(d,m)$, so the crossover should be roughly equivalent to ${\sf COST}_{\sf over}(d,m)$ processors.
The memory cost of $N$ samples of size $d$ is $Nd$, and the memory cost of storing the data is $O(dm)$. The latter may be considerably smaller, and even inconsequential, for neural network models with large number of parameters and smaller input dimension. Nonetheless, we take $O(d(m+N))$ as a proxy for the per process memory requirements of SMC. Meanwhile, if the quantity of interest is known a priori and estimated online, then MCMC samples do not need to be stored, resulting in a memory cost of only $O(dm)$. Of course, if one intends to perform inference in the future without saving samples, then the total memory cost is $O(d(m+NR))$, but the samples could be (thinned and) offloaded at checkpoints, so this can nevertheless be kept under control.
Clearly we require a minimum $N\gg C_{\sf SMC}(d,m)$ in order for the SMC theory to make sense, and we have indeed found that convergence slows or stops if this condition is not satisfied. This can be viewed as sufficient maximum population diversity. Furthermore, the samples must decorrelate sufficiently, leading to a minimum $M$, and they must be be rejuvenated sufficiently often, leading to a minimum $J$. Both of these conditions can also spoil convergence. Conventional advice suggests that it is safe to select $M=J=N=O(\max\{m,d\})$ Chopin et al. [2020]. If $d=m$, then the crossover occurs when $R>d^{2}$, and the memory requirement per processor is also $O(d^{2})$. In practice, we have found that the minimum required $M$ and $J$ often increase much slower than $O(d)$, so the crossover often occurs earlier than $d^{2}$.
Computing and storing the log-likelihood instead of the likelihood itself wherever possible in SMC sampler is preferred in practice to avoid numerical over/underflow. See Appendix A for further details.
4 Simulations
Gaussian m×$\times$×d = 64 Gaussian m×$\times$×d = 1024
m=16 > d m = 8 = d m=4 < d m = 64 > d m = 32 = d m = 16 < d
MSE 1.2332e-07 7.3838e-04 3.1846e+00 6.0035e-07 2.8125e-02 1.9128e+01
No.node 4 2 1 1 20 1
Time 0.9749 1.7513 1.7521 1.3517 6.0139 14.3041
Table 3: The MSE, number of nodes for pSMC, and total time until the crossover point (see Figure 1) for the non-Gaussian example of QST regression. For BNN classification, MSE is replaced with mean KL-divergence.
QST (pCN) BNN (HMC)
(d,m) = (16,6) (d,m) = (64,36) (d,m) = (256,216) iris (d,m) = (15,50) MNIST (d,m) = (1177,50)
MSE/Loss 4.0111e-04 1.5068e-03 4.5424e-05 1.3936e-04 6.7747e-02
No.node 1 1 681 33 26
Time 0.0456 0.6681 29.6914 13.2356 34.5685
4.1 High-dimensional Gaussian cases

### v2

@@ A.1 Pre-conditioned Crank-Nicolson kernel
where the proposal in the original/standard pCN is with $D=I_{d}$. The general version presented above was introduced in [Law, 2014, Cui et al., 2016], and can provide substantially improved mixing when the likelihood informs certain directions much more than others. The scaling matrix $D$ should be chosen according to the information present in the likelihood. A simple and computationally convenient choice which is particularly amenable to use within SMC is to build it from an approximation of the target covariance [Beskos et al., 2018]. In the present work we adopt the simplest and cheapest choice and let $D=\text{diag}(\widehat{\text{var}}[\theta])$. The approximation of the variance, $\widehat{\text{var}}$, will be built from the current population of samples in the SMC case, and adaptively constructed in the MCMC case [Haario et al., 2001, Chen et al., 2016]. See also [Rudolf and Sprungk, 2018, Beskos, 2014, Zahm et al., 2022].
@@
  Inputs: a current state $\theta$, the distribution $\pi_{j}$, a scaling parameter $\beta$.
@@
  Outputs: $\theta$.
A.2 Hamiltonian Monte Carlo
@@
  Inputs: a current state $\theta$, the distribution $\pi_{j}$ and a mass matrix $M_{0}$.
  Generate a initial momentum $q\sim N(0,M_{0})$.
@@
        Generate $\theta_{l\Delta t}$ and $q_{l\Delta t}$ from (LABEL:eqn:leapfrog_step) with $\theta_{0}=\theta$ and $q_{0}=q$
@@
     if $u\leq\min\Big{\{}1,\frac{H(\theta^{\prime},q^{\prime})}{H(\theta,q)}\Big{\}}$ then
        $\theta=\theta^{\prime}$ and $q=q^{\prime}$.
@@
  Outputs: $\theta$.
Hamiltonian Monte Carlo (HMC) kernel [Duane et al., 1987, Neal et al., 2011, Houlsby et al., 2011, Cobb and Jalaian, 2021b] is essentially a gradient-based MCMC kernel on an extended state-space. We first build a Hamiltonian $H(\theta,q)$ with additional auxiliary “momentum” vector $q$ of the same dimension as the state $\theta$
$H(\theta,q)=-\log\pi_{j}(\theta)+\frac{1}{2}q^{\top}M_{0}^{-1}q,$ (11)
where $M_{0}$ is a mass matrix, so that $\frac{1}{Z}\exp(H(\theta,q))$ is a target distribution on the extended space, where the momentum can be simulated exactly $q\sim N(0,M_{0})$.
@@
$\frac{d\theta}{dt}=\frac{\partial H}{\partial q}=M_{0}^{-1}qs,$ (12)
$\frac{dq}{dt}=\frac{\partial H}{\partial\theta}=\nabla_{\theta}\log\pi_{j}(\theta)\,.$
A carefully constructed symplectic-integrator is capable of conserving energy as well, for example the leapfrog integrator [Leimkuhler and Reich, 2004]:
$q_{t+\Delta t/2}=q_{t}+\frac{\Delta t}{2}\frac{dq}{dt}(\theta_{t}),$ (13)
$\theta_{t+\Delta t}=\theta_{t}+\Delta t\frac{d\theta}{dt}(q_{t+\Delta t/2}),$
$q_{t+\Delta t}=q_{t+\Delta t/2}+\frac{\Delta t}{2}\frac{dq}{dt}(\theta_{t+\Delta t}),$
@@
    simulate a random momentum $q\sim N(0,M_{0})$ (hence jumping to a new energy contour);
@@
    approximate the Hamiltonian dynamics using $L$ steps from (LABEL:eqn:leapfrog_step);
@@
    correct numerical error from (ii) with an MH accept/reject step for $\frac{1}{Z}\exp(H(\theta,q))$.
@@
Appendix B Techniques for pSMC in practice
The procedure is defined as follows: for processor $p=1,...,P$,
$Z^{N,p}$ $=\prod_{j=1}^{J}\sum_{i=1}^{N}\omega_{j}^{i,p}=\prod_{j=1}^{J}\Delta K_{j}^{p}\Delta\hat{Z}_{j}^{p}=K^{p}\hat{Z}^{p},$
where $\Delta K_{j}^{p}=\exp(\max_{i}\log(w_{j}^{i,p}))$ and $K^{p}=\prod_{j=1}^{J}\Delta K_{j}^{p}$; $\Delta{\hat{Z}}_{j}^{p}=\sum_{i=1}^{N}\hat{w}_{j}^{i,p}=\sum_{i=1}^{N}\exp(\log(w_{j}^{i,p})-\max_{i}\log(w_{j}^{i,p}))$ and $\hat{Z}^{p}=\prod_{j=1}^{J}\Delta\hat{Z}_{j}^{p}$. For the small number of importance sampling weights $w_{j}^{i,p}$, we work with log-likelihood for unnormalized weights and modify it with a constant $\hat{w}_{j}^{i,p}$.
Using the same trick again on the normalized constant over $P$ processors, we have, for processor $p$,
$Z^{N,p}=K^{p}\hat{Z}^{p}=\exp(\log(\hat{Z}^{p})+\log(K^{p})-\log(K))K,$
where $\log(K)=\max_{p}(\log(\hat{Z}^{p})+\log(K^{p}))$. Since the term $K$ will be cancelled out as we compute the parallel SMC estimator, so we only need the term $\exp(\log(\hat{Z}^{p})+\log(K^{p})-\log(K))$, which could further avoid the computational error due to the small number $K$.
Appendix C Assumptions, Proposition and Lemma
We first present the assumptions.
Assumption C.1.
@@ Appendix C Assumptions, Proposition and Lemma
Assumption C.2.
Let $J\in\mathbb{N}$ be given, for each $j\in\{1,...,J\}$, there exists a $\rho\in(0,1)$ such that for any $(u,v)\in\mathsf{\Theta}^{2}$, $A$ is a measurable set from the space containing all the measurable subset of $\mathsf{\Theta}$,
@@
In order to make use of (6), we require estimates on $f^{N,p}(\zeta)$ both for quantity of interest $\zeta=\varphi$ and $\zeta=1$. We denote that $|\zeta|_{\infty}=\max_{\theta\in\mathsf{\Theta}}|\zeta(\theta)|$ in the following equations. $C_{\zeta}$ denotes the constant depended on the function $\zeta$. Note that using $\varphi$ with one-dimensional output is without loss of generality for our convergence results, and the following proof can be directly generalized to the multi-output function by using the inner product. Proof of the following proposition can be found in Del Moral [2004].
Proposition C.3.
Assume Assumption C.1 and C.2. Then, for any $J\in\mathbb{N}$, there exists a $C>0$ such that for any $N\in\mathbb{N}$, suitable $\zeta:\mathsf{\Theta}\rightarrow\mathbb{R}$,
@@
The following supporting Lemma will be proven in the next section, along with the main theorem 3.1.
Lemma C.4.
@@
Assume Assumption C.1 and C.2. Then, for any $J\in\mathbb{N}$, there is a $C>0$ such that for any $N,P\in\mathbb{N}$, suitable $\zeta:\mathsf{\Theta}\rightarrow\mathbb{R}$,
$\mathbb{E}[(F^{N,P}(\zeta)-f(\zeta))^{2}]\leq\frac{C|\zeta|_{\infty}^{2}}{NP}.$
Appendix D Proofs
D.1 Proof relating to Lemma C.4
@@ Appendix D Proofs
Assume Assumption C.1 and C.2. Then, for any $J\in\mathbb{N}$, there is a $C>0$ such that for any $N,P\in\mathbb{N}$, suitable $\zeta:\mathsf{\Theta}\rightarrow\mathbb{R}$,
$\mathbb{E}[(F^{N,P}(\zeta)-f(\zeta))^{2}]\leq\frac{C|\zeta|_{\infty}^{2}}{NP}.$
Proof.
$\mathbb{E}[(F^{N,P}(\zeta)-f(\zeta))^{2}]$
$=\mathbb{E}\bigg{[}\bigg{(}\frac{1}{P}\sum_{p=1}^{P}(f^{N,r}(\zeta)-f(\zeta))\bigg{)}^{2}\bigg{]}$
$=\frac{1}{P^{2}}\mathbb{E}\bigg{[}\sum_{p=1}^{P}(f^{N,p}(\zeta)-f(\zeta))^{2}+\sum_{p=1}^{P}\sum_{p^{\prime}=1}^{P}(f^{N,p}(\zeta)-f(\zeta))(f^{N,p^{\prime}}(\zeta)-f(\zeta))\bigg{]}$
$=\frac{1}{P^{2}}\sum_{p=1}^{P}\mathbb{E}[(f^{N,p}(\zeta)-f(\zeta))^{2}]+\frac{1}{P^{2}}\sum_{p=1}^{P}\sum_{p^{\prime}=1}^{P}\mathbb{E}[(f^{N,p}(\zeta)-f(\zeta))]\mathbb{E}[(f^{N,p^{\prime}}(\zeta)-f(\zeta))].$ (15)
By Proposition C.3, three expectation terms in (15) are expressed as follows
$\mathbb{E}[(f^{N,p}(\zeta)-f(\zeta))^{2}]\leq\frac{C|\zeta|_{\infty}^{2}}{N},\ \mathbb{E}[(f^{N,p}(\zeta)-f(\zeta))]=0.$
Then, we conclude
$\mathbb{E}[(F^{N,P}(\zeta)-f(\zeta))^{2}]\leq\frac{C|\zeta|_{\infty}^{2}}{NP}.$
D.2 Proof relating to Theorem 3.1
Given Assumptions C.1 and C.2, for suitable values of ($M$,$N$,$J$) there exists a $C_{\varphi}>0$, which depends on $\varphi$, such that for any $P\in\mathbb{N}$,
$\mathbb{E}[(\hat{\varphi}_{\text{pSMC}}-\pi(\varphi))^{2}]\leq\frac{C_{\varphi}}{NP}.$ (16)
Proof.
$\mathbb{E}[(\hat{\varphi}_{\text{pSMC}}-\pi(\varphi))^{2}]$ (17)
$=\mathbb{E}\bigg{[}\bigg{(}\frac{F^{N,P}(\varphi)}{F^{N,P}(1)}-\frac{f(\varphi)}{f(1)}\bigg{)}^{2}\bigg{]}$
$=\mathbb{E}\bigg{[}\bigg{(}\frac{F^{N,P}(\varphi)}{F^{N,P}(1)}-\frac{F^{N,P}(\varphi)}{f(1)}+\frac{F^{N,P}(\varphi)}{f(1)}-\frac{f(\varphi)}{f(1)}\bigg{)}^{2}\bigg{]}$
Applying Cauchy-Schwartz inequality, we have
$\leq 2\mathbb{E}\bigg{[}\bigg{(}\frac{F^{N,P}(\varphi)}{F^{N,P}(1)}-\frac{F^{N,P}(\varphi)}{f(1)}\bigg{)}^{2}\bigg{]}+2\mathbb{E}\bigg{[}\bigg{(}\frac{F^{N,P}(\varphi)}{f(1)}-\frac{f(\varphi)}{f(1)}\bigg{)}^{2}\bigg{]}$
$\leq\frac{2|F^{N,P}(\varphi)|_{\infty}^{2}}{|F^{N,P}(1)|^{2}|f(1)|^{2}}\mathbb{E}[(F^{N,P}(1)-f(1))^{2}]+\frac{2}{|f(1)|^{2}}\mathbb{E}\big{[}(F^{N,P}(\varphi)-f(\varphi))^{2}\big{]}$ (18)
Assume Assumption C.1, there exists a $C_{\varphi}^{\prime}$ such that $\frac{2|F^{N,P}(\varphi)|_{\infty}^{2}}{|F^{N,P}(1)|^{2}|f(1)|^{2}}\leq C_{\varphi}^{\prime}$, and there exists a $C^{\prime}$ such that $\frac{2}{|f(1)|^{2}}\leq C^{\prime}$. Then, following (18), we have
$\mathbb{E}[(\hat{\varphi}_{\text{pSMC}}-\pi(\varphi))^{2}]$ $\leq C_{\varphi}^{\prime}\mathbb{E}[(F^{N,P}(1)-f(1))^{2}]+C^{\prime}\mathbb{E}\big{[}(F^{N,P}(\varphi)-f(\varphi))^{2}\big{]}$
By Lemma C.4 with $\zeta=\varphi$ and $\zeta=1$ respectively, we have
$\leq\frac{C_{\varphi}^{\prime}C}{NP}+\frac{C^{\prime}C|\varphi|_{\infty}^{2}}{NP}$
$\leq\frac{C_{\varphi}}{NP}.$
Appendix E Complementary description of simulations
E.1 Computation of Error bars
Assume running $R$ times of experiments to get $R$ square errors/loss between simulated estimator $\hat{\varphi}$ and the ground truth, SE$(\hat{\varphi})^{r}$ for $r=1,...,R$. Take the MSE as an example, the MSE is the mean of SE$(\hat{\varphi})^{r}$ over $R$ realizations, and the standard error of MSE (s.e.) is computed by
$\frac{\sqrt{\frac{1}{R}\sum_{r=1}^{R}(\text{SE}(\hat{\varphi})^{r}-\text{MSE})^{2}}}{\sqrt{R}}.$ (19)
E.2 Details of Gaussian cases